OpenStudy (anonymous):

the domain and range in a radical expression y=x^2 +2

OpenStudy (markchernioglo):

x = 27

OpenStudy (freckles):

where is the radical exactly?

OpenStudy (freckles):

\[y=\sqrt{x^2+2}\]?

OpenStudy (triciaal):

the range is the set of y values for any value of x then y is positive given y = x^2 + 2 then Y-2 = x^2 x = square root of (Y -2 ) for square root positive and negative values also (y-2) must be positive

Directrix (directrix):

@Ada1959 The vertex of the parabola which is the equation of y= x^2 + 2 is (0 , 2). Because the coefficient of the x^2 term is positive, the parabola opens upward. Therefore the smallest y-value possible is 2. The range is {all y > or = to 2}. There is no reason to exclude any value from the set of possible x-values. The domain is the set of real numbers.